OpenStudy (anonymous):

Given the first term and the common ratio of a geometric sequence, find the first five terms and write the explicit formula. a1 = 1, r = ½

OpenStudy (anonymous):

just like the last one, except this time \(r=\frac{1}{2}\) instead of \(2\)

OpenStudy (anonymous):

first term is 1, second is \(\frac{1}{2}\) third is \(\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\) and so on

OpenStudy (anonymous):

@satellite73 whats the explicit formula, and fourth term is 1/8 and fifth is 1/16

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (texaschic101):

an = a1 * r^n-1 a1 is the first term r is the common difference n is the number of the term you want to find.

OpenStudy (anonymous):

ty!!!!!!!!